OpenStudy (anonymous):

Evaluate and write the solution in scientific notation: 4.2x10^5 ----------- 2.1x10^2

OpenStudy (anonymous):

answers: 2 x 107 2 x 103 2 x 102.5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge\text{Guide the asker to a solution}$$

OpenStudy (anonymous):

could you throw in some steps that I should do?

OpenStudy (skullpatrol):

$$\Huge \frac{4.2*10^5}{2.1*10^2}$$

OpenStudy (skullpatrol):

You need to do division...

OpenStudy (anonymous):

0.000042 is the answer to bottom one?

OpenStudy (anonymous):

i mean top one*

OpenStudy (skullpatrol):

$$\Huge \frac{ac}{bd} =\frac{a}{b} * \frac{c}{d}$$

OpenStudy (skullpatrol):

$$\Huge \frac{4.2*10^5}{2.1*10^2} =?$$

OpenStudy (anonymous):

420000 ------ 210

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

OK that works... now what?

OpenStudy (skullpatrol):

?

OpenStudy (anonymous):

the next step? I don't know ;/

OpenStudy (skullpatrol):

Divide 420,000 by 210.

OpenStudy (skullpatrol):